OpenStudy (anonymous):

Calculus I- The equation of the tangent to the curve x^2=4y at the point on the curve where x=-2 is...

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

and you thoughts about this are?

OpenStudy (amistre64):

to create a line, we can use a point and a slope. how do we define these parts?

OpenStudy (anonymous):

y=x^2/4

OpenStudy (anonymous):

point is (-2, 1) i believe

OpenStudy (amistre64):

good, and the slope?

OpenStudy (anonymous):

x/2 is the slope

OpenStudy (amistre64):

m = x/2, for a given x yes

OpenStudy (anonymous):

so how do we find the tangent of the curve at this point? is it just y=-1?

OpenStudy (amistre64):

the slope is -1

OpenStudy (amistre64):

what line equation formats do we know of?

OpenStudy (anonymous):

uh....y=mx+b

OpenStudy (amistre64):

y = mx + b is good, but only if we know a specific point, the y intercept lets use point slope format, since we have a general point and a slope y - p1 = m(x-px)

OpenStudy (amistre64):

ugh ... not enough mt dew yet y - py = m(x-px)

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

is that an: ok it makes sense? or an: ok now what do i do?

OpenStudy (anonymous):

ok now what do i do?

OpenStudy (amistre64):

plug in the values that we found m = -1 the point (px,py) = (-2,1)

OpenStudy (anonymous):

y-1=-1(x+2)

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

y=-x-1

OpenStudy (anonymous):

so x+y+1=0

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

as a dbl chk http://www.wolframalpha.com/input/?i=y%3Dx%5E2%2F4%2C+y%3D-%28x%2B2%29%2B1 youre welcome :)